HUBBLE SPACE TELESCOPE DAILY REPORT #1939 PERIOD COVERED: 0000Z (UTC) 08/11/97 - 0000Z (UTC) 08/012/97 Daily Status Report as of 224/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Seven Sets of WF/PC-2 7620 (WF/PC-2 CYCLE 7 Standard Darks) The WF/PC-2 was used to make a series of Cycle 7 standard dark calibration images. This dark calibration program obtains dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed WF/PC-2 7626 (UV Throughput) The WF/PC-2 was used to make observations of PPL-15 as part of as ultraviolet throughput monitoring program. The observations were completed as planned, and no problems were reported. 1.3 Completed STIS/CCD 7659 (Daily Darks to Update Acquisition Bad Pixel Table) The Space Telescope Imaging Spectrograph (CCD) was used to make dark calibration images in order to update the acquisition bad pixel table. Daily hot pixel lists can also be used for science data analysis. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed Two Sets of WF/PC-2 7621 (WF/PC-2 Cycle 7 Supplemental Darks PT1/3) The WF/PC-2 was used to make supplemental dark calibration images. This set of dark calibrations is intended to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed NIC/1/2/3 7775 (NICMOS Earth Flats) The Near Infrared Spectrograph Multi-Object Camera (NICMOS) was used to make Earth flat calibration observations. This proposal obtains a series of Earth flats for a subset of narrow-, medium-and broad-band NIC 1, 2 and 3 filters. These flats will be used to generate the master flats for narrow-band filters and to determine the OTA illumination pattern as a function of wavelength and camera. The observations were completed as scheduled, and no problems were noted. 1.6 Completed NIC/1/2/3/STIS/CCD 7267 (Infrared Imaging and Polarimetry Observations of the Obscured Nucleus of Centaurus A) The Near Infrared Camera (Cameras 1, 2 and 3) and the Space Telescope Imaging Spectrograph (CCD) was used to make infrared and polarimetry observations of the inner nuclear region of Centaurus A. This project complements and extends the studies we performed with HST in the optical bands and in the infrared with ground based data. The observations were completed as scheduled, and no anomalies were reported. 1.7 Completed NIC/2 7331 (Near-IR Properties of the Bulges of Spiral Galaxies) The Near Infrared Camera (Camera Two) was used to make infrared observations of the bulge in spiral galaxy ESO572G22. Near-IR data at HST resolution are essential for correcting for the effect of dust and will enable us to derive robust nuclear properties for the sample. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed STIS/CCD 7635 (STIS CCD Hot Pixel Annealing) The Space Telescope Imaging Spectrograph (CCD) was used to test the hot pixel annealing procedure. The effectiveness of the CCD hot pixel annealing process is assessed by measuring the dark current behavior before and after annealing and by searching for any window contamination effects. The observations were completed as planned, and no problems were reported. 1.9 Completed NIC 2 7498 (NICMOS Imaging Survey of Distant Radio Galaxies. Ellipticals at Z > 2 ?) The Near Infrared Camera (Camera 2) was used to make high resolution and high signal-to-noise ratio imaging survey observations of the radio galaxy MRC0406-24. The existence and photometric properties of massive elliptical galaxies at high redshift provide a powerful probe of galaxy evolution. The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed Three Sets of WF/PC-2 6802 (Cosmology With the Medium Deep Survey) The WF/PC-2 was used to make observations taken in parallel with the prime science observations as part of the Medium Deep Sky Survey. The observations were executed as scheduled, and no problems were reported. 1.11 Completed WF/PC-2 6802 (Cosmology With the Medium Deep Survey) The WF/PC-2 was used to make observations taken in parallel with the prime science observations as part of the Medium Deep Sky Survey. The observations were executed as scheduled, and no problems were reported. 1.12 Completed S/C 5582 (FOC/48 Monthly Activation (Camera Section Only): Cycle 4) This proposal was executed as scheduled, and no anomalies were noted. 1.13 Completed WF/PC-2 6587 (Black Holes and Cores of Early-Type Galaxies) The WF/PC-2 was used to make observations of early-type galaxy NGC-1016 in order to study its nuclear core region and its possible black hole. The campaign to find black holes in the centers of galaxies using HST has so far been highly successful. The observations were completed as planned, and no problems were reported. 1.14 Completed NIC/1/2/3 7447 (The Cosmological Evolution of Quasar Host Galaxies) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of galaxy PKS0440-00 as part of a survey to study the cosmological evolution of galaxies which host quasars. Other point spread function observations were taken to test the calibration of the cameras. The observations were executed as scheduled, and no anomalies were noted. 1.15 Completed FGS 7682 (FGS #3 Blue Transfer Function Reference Search Calibration) The Fine Guidance Sensor #3 was used to make observations of candidate star HD38666 in order to perform the blue Transfer Function reference search calibration. This proposal will measure the transfer function of a number of blue candidate stars in order to verify a single star which can be used for the blue calibration of the Trans Mode. The observations were exected as scheduled, and no anomalies were reported. 1.16 Completed NIC/1/2/3 7726 (NICMOS One Orbit Parallel) The Near Infrared Camera (Cameras One, Two and Three) was used to make a one orbit parallel calibration observations. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 10 Successful : 10 Scheduled Reacquisitions : 6 Successful : 6 2.2 FHST Updates: Scheduled : 17 Successful : 17 2.3 Operations Notes: No operations anomalies were noted. Nominal Operations. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH